OpenStudy (anonymous):

What is the domain of sqrt(1-x^2) ?

OpenStudy (anonymous):

Well, that depends. :D

OpenStudy (anonymous):

Are you referring to a complex worth function?

OpenStudy (anonymous):

ok the question assigned is what is the domain of f(x,y) = sqrt(1-x^2) - sqrt(1-y^2)

OpenStudy (anonymous):

I'm a bit rusty on this stuff

OpenStudy (anonymous):

all reals where 1-(x^2) >= 0 because you can't have a nevgative in the square root . simplified to 1 >= x^2 you get x > or = +- 1

OpenStudy (anonymous):

You can have negative in the root.

OpenStudy (anonymous):

Simply use the definition that i=sqrt(-1). (Which is why you need to know if the function is complex/real.

OpenStudy (anonymous):

Oh gotcha, It's a real function not complex

OpenStudy (anonymous):

Then just state what is required; 1-x^2>=0 and 1-y^2>=0

OpenStudy (anonymous):

Do you know where to go from here?

OpenStudy (anonymous):

Ya I do, that was a useful response, thanks

OpenStudy (anonymous):

No problem man, I can tell you're here to learn! :) Just don't forget that sqrt(x^2) is generally not the same thing as x.

OpenStudy (anonymous):

But instead sqrt(x^2)=|x|

OpenStudy (anonymous):

So i was right.. i know what imaginary numbers are, Noliec. thanks

OpenStudy (anonymous):

cool, thanks

OpenStudy (anonymous):

You were breaking the Code of Conduct on openstudy quite straight forward, footballa.